OpenStudy (anonymous):

A circle has a radius of 12 centimeters. If 6/7 of the circle is cut away, what is the area of the remaining portion?

Parth (parthkohli):

It is obvious that 1/7 of the area is remaining. First calculate the area then multiply that by 1/7.

OpenStudy (anonymous):

51.43 am i going right?

OpenStudy (anonymous):

A = pi*r^2 = 452.4 (1dp ) 1/7 of the area remaining so (1/7) * 452.4 = 64.6 (1dp) cm^2

OpenStudy (anonymous):

oh

Parth (parthkohli):

\( \color{Black}{\Rightarrow\Large {3.14 \times 144 \over 7} }\)

OpenStudy (anonymous):

yeah thanks